# Towards Resilient Safety-driven Unlearning for Diffusion Models against Downstream Fine-tuning

### v1

@@ 4.1 Experimental Setup
Baselines & Diffusion Models. We evaluate our method against 6 text-to-image diffusion models, including the Stable Diffusion v1.4 rombach2022ldm and 5 state-of-the-art baselines for unsafe concept erasure: ESD gandikota2023erasing , SafeGen li2024safegen , AdvUnlearn zhang2024defensive , and two variants of LCFDSD pan2024leveraging (i.e., LCFDSD-NG and LCFDSD-LT). We directly use their provided checkpoints or run their official code with default hyperparameters to obtain their unlearned models. Note that the baselines are mostly implemented and evaluated based on SD v1.4 and mainly focus on the unsafe concept of sexual, and most of them do not provide extensions to other types of unsafe content or to other model architectures. Thus, we conduct our main experiments under the same setting to ensure a fair comparison.
@@
Fine-tuning Methods & Datasets. Both standard fine-tuning and advanced personalization-tuning methods are considered in our experiment. For standard fine-tuning, we adopt (a filtered subset of) two datasets: DreamBench++ peng2025dreambench and DiffusionDB wang2023diffusiondb , which contain high-quality images of human characters, artistic styles, etc. We use these datasets to serve as a representative for benign fine-tuning. In addition, we follow pan2024leveraging and use the Harmful-Imgs dataset, which consists of a diverse set of sexually explicit prompt-image pairs, to understand the resilience of ResAlign when the fine-tuning dataset (partially) contains inappropriate data. Beyond these, we further evaluate the generalizability of our method under more advanced personalization-tuning methods, including LoRA hu2022lora , DreamBooth ruiz2023dreambooth , SVDiff han2023svdiff , and CustomDiffusion kumari2023multi , with more personalization datasets including Pokémon reach_vb_pokemon_blip_captions , Dog peng2025dreambench , ArtBench liao2022artbench , and VGGFace2-HQ chen2023simswap . Finally, we employ the I2P schramowski2023sld and Unsafe qu2023unsafe datasets for measuring harmful generation capabilities, and the COCO chen2015microsoft dataset for assessing benign generation capability.
@@
Evaluation Metrics. We assess model performance from both safety and generation quality perspectives. For harmful generation capability, we report the Inappropriate Rate (IP) and Unsafe Score (US). IP is computed by generating images using the I2P dataset schramowski2023sld prompts and measuring the proportion of outputs flagged as harmful by either the NudeNet nudenetclassifier or Q16 schramowski2022can detector. Similarly, US is computed on images generated from the Unsafe dataset qu2023unsafe prompts with another pretrained NSFW classifier (i.e., MHSC qu2023unsafe ). A lower IP ($\downarrow$) and US ($\downarrow$) indicate weaker unsafe tendencies, meaning the model poses less risk of misuse and is less likely to expose benign users to inappropriate content li2024art ; schramowski2023sld . Note that, different from Pan et al. pan2024leveraging who evaluate model safety on the full set of unsafe prompts, we follow previous works li2024safegen ; zhang2024defensive and mainly focus on the sexual category in both datasets to ensure a more targeted and fair comparison. For benign generation, we use three standard metrics: FID heusel2017gans , CLIP score radford2021learning , and Aesthetics Score laion_ai_aesthetic_predictor , which are widely used by previous works pan2024leveraging ; zhang2024defensive , to measure the ability of the unlearned models in generating benign concepts. Finally, we also follow previous work han2023svdiff ; ruiz2023dreambooth to use CLIP-I and CLIP-T ruiz2023dreambooth , as well as DINO score (oquab2024dinov, ) to measure the performance of our model in generating personalized concepts after fine-tuning.
@@
Before Fine-tuning Fine-tuned on DreamBench++ Fine-tuned on DiffusionDB FID ↓ / CLIP Score ↑ / Aesthetics Score ↑
@@
Implementation Details. By default, we initialize our model using SD v1.4 and train it following Alg. 2. For our meta-learning, the distribution of configurations $\pi(\mathcal{C})$ include the learning rates of $[1\times 10^{-4},1\times 10^{-5},1\times 10^{-6}]$, the step of $[5,10,20,30]$, the fine-tuning loss (i.e., $\mathcal{L}_{\text{FT}}$) of both standard denoising loss and the prior-preserved denoising loss ruiz2023dreambooth , the algorithm of both full-parameter fine-tuning and LoRA hu2022lora , and the optimizer of both SGD and Adam. For each adaptation simulation, we select a subset of 50 images from the DiffusionDB wang2023diffusiondb dataset to act as the dataset (i.e., $\mathcal{D}_{\text{FT}}$) and a random combination of the above configurations to form $\mathcal{C}$. The outer loop learning rate and steps are set to $2\times 10^{-4}$ and 120, respectively, and the hyperparameters are empirically set as $\alpha=1,\beta=0.8,K=5,\gamma=1$. The full training process is done on a single NVIDIA RTX A100 GPU with a memory consumption of $\sim 56$GB, and the training time is approximately $58$ GPU minutes. In our evaluation, all fine-tuning is full-parameter fine-tuning on the respective dataset for 200 steps with a batch size of $1$ and a learning rate of $1\times 10^{-5}$ by default. To mitigate the randomness during training and evaluation, are experiments are independently repeated 3 times with different random seeds, and we report the averaged results unless otherwise stated.
@@ 4.2 Experimental Results
Fig. 1: Visualization results on harmful generation. Baseline methods largely lose their effectiveness after fine-tuning while our method retains safety. The black blocks are added by authors to avoid disturbing readers.
As shown in Tab. 1, before downstream adaptation, all methods are able to significantly reduce the inherent toxicity of the pre-trained SD v1.4 model, as reflected by their lower IP and US scores. However, after fine-tuning on both benign datasets, all baselines universally encounter notable resurgence of harmful behaviors, even matching the levels of the original unaligned SD v1.4 (e.g., AdvUnlearn on DiffusionDB and SafeGen on both datasets). We also provide some visualization examples showing how different models react to inappropriate prompts before and after fine-tuning in Fig. 1. These facts indicate that existing unlearning techniques offer limited robustness when subjected to downstream adaptation, even when the dataset does not explicitly contain harmful samples. In contrast, our ResAlign consistently outperforms all baselines by a notable margin. Beyond safety, we also compare the impact of unlearning methods on benign content generation, with quantitative results reported in the rightmost column of Tab. 1 and qualitative results in Fig. 3. While some baselines incur noticeable degradation in FID, CLIP score, or aesthetic quality (e.g., LCFDSD), our method introduces minimal performance drop and remains competitive with the original SD v1.4 and most baselines. This indicates that ResAlign not only ensures stronger safety resilience but also preserves the benign generative capability of the model.
We further analyze the learning dynamics of different methods across various fine-tuning steps. As can be seen in Fig. 2, baseline methods show notable fluctuations and a rapid increase in IP with only
@@
a small number of fine-tuning steps, suggesting that their unlearning is brittle in the weight space. ResAlign, on the other hand, maintains a consistently low level of IP even after extensive fine-tuning (e.g., up to 500 steps), with remarkably smaller fluctuations. These results confirm that our method not only effectively suppresses harmful behavior effectively but also endows the model to better withstand downstream re-acquisition of unsafe capabilities.
@@
DreamBooth ruiz2023dreambooth + Dog ruiz2023dreambooth SD v1.4 0.3646 0.1884 0.2671 0.9532 0.8763
Ours 0.0021 0.0260 0.2631 0.9466 0.8605
@@
LoRA hu2022lora + Pokémon reach_vb_pokemon_blip_captions SD v1.4 0.3269 0.1977 0.3178 0.6544 0.4549
Ours 0.0032 0.0322 0.3175 0.6512 0.4332
@@
As shown in Tab. 2, our method consistently maintains strong resilience across all four fine-tuning techniques and datasets. Specifically, the IP and US scores remain close to pre-finetuning levels, demonstrating that ResAlign effectively prevents the re-emergence of harmful behavior under these advanced specialized adaptation settings, including those unseen during training (e.g., SVDiff and CustomDiffusion). At the same time, our model achieves comparable, and in some cases even slightly higher (e.g., on ArtBench), CLIP-I, CLIP-T, and DINO scores compared to the original SD v1.4. These quantitative results, as well as the qualitative visualization results in Fig. 3 suggest that our method well-preserved the personalization capability of the unlearned model, i.e., our ResAlign will not interfere with benign downstream tasks.
Effectiveness across More Fine-tuning Configurations. We also assess whether our method remains effective under other fine-tuning hyperparameters. Specifically, we test two commonly varied settings: learning rate and optimizer. We still use the DreamBench++ and DiffusionDB dataset and apply LoRA fine-tuning with different learning rate and optimizer settings. As shown in Tab. 4 and 5, ResAlign continues to perform well across these configurations, maintaining low harmfulness across different settings. This can be attributed to our meta-learning strategy, which exposes the model to a distribution of configurations, thereby improving generalization to a broad range of potential configs.
Effectiveness on Contaminated Data. Lastly, we consider a challenging scenario where the fine-tuning data is either intentionally or inadvertently contaminated with unsafe content. To evaluate this, we mix our datasets with the Harmful-Imgs dataset pan2024leveraging at varying contamination ratios, and use this contaminated data to fine-tune the unlearned models. As shown in Fig. 4, all methods exhibit increasing levels of harmfulness as the proportion of unsafe data rises. Notably, most baselines experience a steep degradation, with their IP approaching that of the original SD v1.4 when 20% or more of the data is unsafe. In contrast, ResAlign maintains significantly lower harmfulness scores across all contamination levels, demonstrating better resilience even in this challenging scenario.
Overall, we argue perfect resilience against harmful data is inherently difficult—if not impossible—as, an adversary can, afterall, treat the recovery of unsafe behavior as a new learning task deng2024sophon , which pre-trained diffusion models are adept at due to their strong generalization and few-shot
@@
Tab. 3: Effect of components.
Component Metric
Hyper. Meta. (D) Meta. (C) IP $\downarrow$ FID $\downarrow$
-- -- -- 0.2266 18.24
$\checkmark$ -- -- 0.1826 18.07
$\checkmark$ $\checkmark$ -- 0.0322 18.35
$\checkmark$ $\checkmark$ $\checkmark$ 0.0186 18.18
Fig. 5: Effect of $\gamma$.
Ablation Study. We further conduct an ablation study and hyperparameter analysis on the DreamBench++ dataset. From Tab. 4.2, we can see that every component contributes to our ResAlign, with hypergradient approximation (Hyper.) enabling reducing regained harmfulness and meta-learning on dataset (Meta. (D)) and configurations (Meta. (C)) notably mitigate overfitting and enhance generalization. Besides, the hyperparameter $\gamma$ controls the degree of proximity to the base model. A larger $\gamma$ can better approximate the ground-truth fine-tuned model, yet also leads to higher instability due to larger variance in hypergradient approximation, as shown in Fig. 5. Overall, our method is stable when $\gamma$ is in a reasonable range (i.e., $0\sim 1$). Thus, a wide range of $\gamma$ can be selected for ResAlign.

### v2

@@ 4.1 Experimental Setup
Baselines & Diffusion Models. We compare our method with 6 text-to-image diffusion models, including SD v1.4 rombach2022ldm, ESD gandikota2023erasing, SafeGen li2024safegen, AdvUnlearn zhang2024defensive, and two variants of LCFDSD pan2024leveraging (i.e., LCFDSD-NG and LCFDSD-LT). We directly use their provided checkpoints or run their official code with default hyperparameters to obtain their unlearned models (see more details in Appendix B.1). Note that the baselines are mostly implemented based on SD v1.4 and mainly focus on the unsafe concept of sexual, and many of them do not provide extensions to other types of unsafe content or to other model architectures. Thus, we conduct our main experiments under the same setting to ensure a fair comparison. We also validate ResAlign on other diffusion models in Sec. 4.2.
@@
Fine-tuning Methods & Datasets. Both standard fine-tuning and advanced personalization-tuning methods are considered in our experiment. For standard fine-tuning, we adopt a selected subset of two datasets: DreamBench++ peng2025dreambench and DiffusionDB wang2023diffusiondb, which contain high-quality images of human characters, artistic styles, etc. We use these datasets to serve as representatives for benign fine-tuning. In addition, we follow pan2024leveraging and use the Harmful-Imgs dataset, which consists of a diverse set of sexually explicit prompt-image pairs, to understand the resilience of ResAlign when the fine-tuning dataset (partially) contains inappropriate data. Beyond these, we further evaluate the generalizability of our method under more advanced personalization-tuning methods, including LoRA hu2022lora, DreamBooth ruiz2023dreambooth, SVDiff han2023svdiff, and CustomDiffusion kumari2023multi, with more personalization datasets including Pokémon reach_vb_pokemon_blip_captions, Dog peng2025dreambench, ArtBench liao2022artbench, and VGGFace2-HQ chen2023simswap. Finally, we employ the I2P schramowski2023sld and Unsafe qu2023unsafe datasets for measuring harmful generation capabilities, and the COCO chen2015microsoft dataset for assessing benign generation capability. More details are in Appendix B.2.
@@
Evaluation Metrics. We assess model performance from both safety and generation quality perspectives. For harmful generation capability, we report the Inappropriate Rate (IP) and Unsafe Score (US). IP is computed by generating images from I2P dataset prompts schramowski2023sld and measuring the proportion of outputs identified as harmful, where sexual content is detected by the NudeNet detector nudenetclassifier. Similarly, US is computed on images generated from the Unsafe dataset qu2023unsafe prompts with another pretrained NSFW classifier (i.e., MHSC qu2023unsafe). A lower IP ($\downarrow$) and US ($\downarrow$) indicate weaker unsafe tendencies, meaning the model poses less risk of misuse and is less likely to expose benign users to inappropriate content li2024art; schramowski2023sld. Note that, unlike Pan et al. pan2024leveraging who evaluate model safety on the full set of unsafe prompts, we focus on the sexual category subset in both datasets to ensure a more targeted and fair comparison. For benign generation, we use FID heusel2017gans, CLIP score radford2021learning, and Aesthetics Score laion_ai_aesthetic_predictor, which are widely used by previous works pan2024leveraging; zhang2024defensive, to measure the ability of the unlearned models to generate benign concepts. Finally, we also follow previous work han2023svdiff; ruiz2023dreambooth to use CLIP-I and CLIP-T ruiz2023dreambooth, as well as DINO score (oquab2024dinov) to measure the performance of our model in generating personalized concepts after fine-tuning. More details are presented in Appendix B.3.
@@
Before
Fine-tuning
Fine-tuned on
DreamBench++
Fine-tuned on
DiffusionDB
FID ↓ / CLIP Score ↑
/ Aesthetics Score ↑
@@
Implementation Details. By default, we initialize our model using SD v1.4 and train it following Alg. 2. For our meta-learning, the distribution of configurations $\pi(\mathcal{C})$ includes the learning rates of $1\times 10^{-4},1\times 10^{-5},1\times 10^{-6}]$, the steps of $[5,10,20,30]$, the fine-tuning loss (i.e., $\mathcal{L}_{\text{FT}}$) of both standard denoising loss and the prior-preserved denoising loss ruiz2023dreambooth, the algorithm of both full-parameter fine-tuning and LoRA hu2022lora, and the optimizer of both SGD and Adam. For each adaptation simulation, we select a random combination of the above configurations to form $\mathcal{C}$. Some abnormal configurations (e.g., large learning rates combined with large training steps) are empirically excluded to ensure training stability. The outer loop learning rate is set to $2\times 10^{-4}$. Training is performed on a single NVIDIA RTX A100 GPU until convergence, which typically requires $\sim 1$ GPU hour. During training, the peak and average memory consumption are $\sim 56$ GB and $\sim 24$ GB, respectively. In our evaluation, all fine-tuning is full-parameter fine-tuning on the respective dataset for 200 steps with a batch size of $1$ and a learning rate of $1\times 10^{-5}$ by default (more details in Appendix B.4). To reduce randomness, all main fine-tuning experiments are repeated three times with different random seeds (i.e., three independent runs), and we report the averaged results. Besides, we also report the standard error of the mean in our figures as the error area to help understand the statistical significance of our results. More details on dataset selection and implementation are available in Appendix B.
@@ 4.2 Experimental Results
Fig. 1: Visualization of harmful generation. Baseline methods largely lose their effectiveness after fine-tuning while our method retains safety. The black blocks are added by the authors to avoid disturbing readers.
As shown in Tab. 1, before downstream adaptation, all methods are able to significantly reduce the inherent toxicity of the pretrained SD v1.4 model, as reflected by their lower IP and US scores. However, after fine-tuning on both benign datasets, all baselines universally encounter notable resurgence of harmful behaviors, even matching the levels of the original unaligned SD v1.4 (e.g., AdvUnlearn on DiffusionDB and SafeGen on both datasets). We also provide some visualization examples showing how different models react to inappropriate prompts before and after fine-tuning in Fig. 1. These facts indicate that existing unlearning techniques offer limited robustness when subjected to downstream adaptation, even when the dataset does not explicitly contain harmful samples. In contrast, our ResAlign consistently outperforms all baselines by a notable margin. Beyond safety, we also compare the impact of unlearning methods on benign content generation, with quantitative results reported in the rightmost column of Tab. 1 and qualitative results in Fig. 3. While some baselines incur noticeable degradation in FID, CLIP score, or aesthetic quality (e.g., LCFDSD), our method introduces minimal performance drop and remains competitive with the original SD v1.4 and most baselines. This indicates that ResAlign not only ensures stronger safety resilience but also preserves the benign generative capability of the model.
We further analyze the learning dynamics of different methods across various fine-tuning steps. As
@@
can be seen in Fig. 2, baseline methods show notable fluctuations and a rapid increase in IP with only a small number of fine-tuning steps, suggesting that their unlearning is brittle in the weight space. ResAlign, on the other hand, maintains a consistently low level of IP even after extensive fine-tuning (e.g., up to 500 steps), with remarkably smaller fluctuations. These results collectively confirm that our method not only suppresses harmful behavior effectively but also allows the unlearned model to better withstand re-acquisition of unsafe capabilities under realistic downstream fine-tuning steps.
@@
DreamBooth ruiz2023dreambooth + Dog ruiz2023dreambooth
SD v1.4 0.3645 0.1883 0.2671 0.9532 0.8763
Ours 0.0021 0.0000 0.2631 0.9466 0.8605
@@
LoRA hu2022lora + Pokémon reach_vb_pokemon_blip_captions SD v1.4 0.3269 0.1683 0.3178 0.6544 0.4549
Ours 0.0032 0.0033 0.3175 0.6512 0.4332
@@
As shown in Tab. 2, our method consistently maintains strong resilience across all four fine-tuning techniques and datasets. Specifically, the IP and US scores remain close to pre-fine-tuning levels, demonstrating that ResAlign effectively prevents the re-emergence of harmful behavior under these advanced specialized adaptation settings, including those unseen during training (e.g., SVDiff and CustomDiffusion). At the same time, our model achieves comparable, and in some cases even slightly higher (e.g., on ArtBench), CLIP-I, CLIP-T, and DINO scores compared to the original SD v1.4. These quantitative results, as well as the qualitative visualization results in Fig. 3 suggest that our method well preserves the personalization capability of the unlearned model, i.e., ResAlign does not interfere with benign downstream tasks.
Effectiveness across More Fine-tuning Configurations. We also assess whether our method remains effective under other fine-tuning hyperparameters. Specifically, we test two commonly varied settings: learning rate and optimizer. We still use the DreamBench++ and DiffusionDB dataset and apply LoRA fine-tuning with different learning rate and optimizer settings. As shown in Tab. 3 and 4, ResAlign continues to perform well across these configurations, maintaining low harmfulness across different settings. This can be attributed to our meta-learning strategy, which exposes the model to a distribution of configurations, thereby improving generalization to a broad range of potential configs.
Effectiveness on Contaminated Data. We further consider a challenging scenario where the fine-tuning data is either intentionally or inadvertently contaminated with unsafe content. To evaluate this, we mix our datasets with the Harmful-Imgs dataset pan2024leveraging at varying contamination ratios, and use this contaminated data to fine-tune the unlearned models. As shown in Fig. 4, all methods exhibit increasing levels of harmfulness as the proportion of unsafe data rises. Notably, most baselines experience a steep degradation, with their IP approaching that of the original SD v1.4 when 20% or more of the data is unsafe. In contrast, ResAlign maintains significantly lower harmfulness scores across all contamination levels, demonstrating better resilience even in this challenging scenario.
Overall, we argue perfect resilience against harmful data is inherently difficult-if not impossible-as, an adversary can, after all, treat the recovery of unsafe behavior as a new learning task deng2024sophon, which pretrained diffusion models are adept at due to their strong generalization and few-shot
@@
Generalizability across More Diffusion Models. By design, our method is model-agnostic and can be readily applied to diverse diffusion architectures. To further verify the applicability of ResAlign
Tab. 5: Effectiveness of ResAlign across more diffusion models. The metric is IP.
Model No FT DB++ Diff.DB
SD v2.0 0.004 0.031 0.078
SDXL 0.033 0.044 0.059
AnythingXL 0.015 0.062 0.087
PonyDiffusion 0.023 0.045 0.067
across a wider range of models, we evaluate it on both official diffusion models (Stable Diffusion v2.0 stabilityai_stable_diffusion_2 and SDXL podell2024sdxl) and two widely adopted community variants (AnythingXL anything_xl_90854 and PonyDiffusion pony_diffusion_v6_xl_297064). Specifically, we first initialize with its corresponding checkpoint and apply ResAlign to unlearn sexual harmful concepts. We then evaluate the safety of each unlearned model (No FT) and subsequently perform LoRA fine-tuning on DreamBench++ and DiffusionDB to assess its safety after downstream fine-tuning. As summarized in Tab. 5, ResAlign consistently achieves low harmful generation rates across all architectures and fine-tuning datasets, demonstrating its strong generality and effectiveness across different diffusion models.
Tab. 6: Effect of components.
Component Metric
Hyper. Meta. (D) Meta. (C) IP $\downarrow$ FID $\downarrow$
- - - 0.2266 18.24
$\checkmark$ - - 0.1826 18.07
$\checkmark$ $\checkmark$ - 0.0322 18.35
$\checkmark$ $\checkmark$ $\checkmark$ 0.0186 18.18
Fig. 5: Effect of $\gamma$.
Ablation Study. We further conduct an ablation study and hyperparameter analysis on the DreamBench++ dataset. From Tab. 4.2, we can see that every component contributes to our ResAlign, with hypergradient approximation (Hyper.) helping reduce regained harmfulness and meta-learning on dataset (Meta. (D)) and configurations (Meta. (C)) notably mitigates overfitting and enhance generalization.
Hyperparameter Analyses. Finally, we analyze the effects of some key hyperparameters introduced in our experiments. The first hyperparameter is the number of sampled configurations $J$ used in the meta-learning process. As shown in Tab. 7, the overall safety after fine-tuning improves as the number of sampled configurations increases. This aligns with intuition, since more meta-learned configurations provide a better estimate of the hypergradient and avoids overfitting, albeit at the cost of longer training time. Interestingly, we observe that increasing $J$ from 1 to 3 provides a notable boost in IP, while further increasing it to 5 yields only marginal additional gains, possibly because $J=3$ already provides sufficiently generalizable hypergradients. Besides,
Tab. 7: Effect of the number of sampled configurations $J$. The metric is IP after fine-tuned on DiffusionDB.
$J$ 1 3 5
Performance (IP) 0.092 0.068 0.054
Training time (min) 43.5 58.2 73.9
$\gamma$ controls the degree of proximity to the base model. To study its impact, we train ResAlign using different $\gamma$ and evaluate the corresponding model’s IP after being fine-tuned on DreamBench++. A larger $\gamma$ can better approximate the ground-truth fine-tuned model, yet also leads to higher instability due to larger variance in hypergradient approximation, as shown in Fig. 5. Overall, our method is stable when $\gamma$ is in a reasonable range (i.e., $0\sim 1$). Thus, a relatively wide range of $\gamma$ can be selected for ResAlign.
5 Discussion
Comparison with More Baselines. In addition to the baselines discussed in Tab. 1, we further compare ResAlign with two additional relevant methods. The first is IMMA zheng2024imma, which employs a bi-level optimization framework to prevent the model from learning certain concepts. Although
Tab. 8: Comparison with other related & concurrent works. The metric is IP after fine-tuning.
Method DB++ Diff.DB
IMMA zheng2024imma 0.0752 0.1321
Meta-Unlearning gao2024meta 0.1128 0.1515
Ours 0.0186 0.0687
originally designed to prevent unauthorized personalized learning, it can also be adapted to our safety-driven unlearning setting for evaluation. The second is the concurrent work Meta-Unlearning gao2024meta, which introduces a meta-objective that penalizes the norm of harmful gradients and encourages conflicts between harmful and benign gradients, thereby slowing down relearning harmful concepts while also encouraging the model to “deconstruct” its benign utility when fine-tuned on harmful data. As shown in Tab. 8, while both methods achieve better results than the baselines in Tab. 1, they remain less effective than ResAlign. We attribute this to two main factors. First, both IMMA and Meta-Unlearning adopt a first-order approximation, which essentially treats the hypergradient’s Jacobian as an identity matrix. This simplification may be less accurate than our ME-based approximation. Second, both methods use fixed datasets and configuration when simulating fine-tuning, which may be less generalizable than our cross-configuration meta learning design. We provide further discussions on more related work in Appendix D and empirical comparisons with additional baselines in Appendix C.2.
Alternative Design. In our main experiments, we adopt the negative denoising score matching loss as $\mathcal{L}_{\text{harmful}}$ (i.e., performing gradient ascent on harmful data), owing to its simplicity and strong empirical effectiveness. However, as observed in prior studies zhang2025understanding; zheng2024imma; li2025forget and our experiments (e.g., Fig. 1), this GA-style loss may cause the model to generate distorted, mosaic-like output images for harmful or semantically related prompts, which may negatively affect user experience to some extent.
Tab. 9: Effectiveness on ResAlign when combined with unlearning loss from ESD gandikota2023erasing and CA kumari2023ablating.
Method No FT DB++ Diff.DB
ResAlign + ESD 0.0610 0.0820 0.1182
ResAlign + CA 0.1230 0.1611 0.1665
Fortunately, ResAlign is a loss-agnostic framework. Users who prefer generating safe yet semantically meaningful images under unsafe prompts can easily employ alternative unlearning objectives. To validate this hypothesis, we replace $\mathcal{L}_{\text{harmful}}$ with the unlearning losses from ESD and CA to train the corresponding unlearned models, and evaluate their safety both immediately after unlearning (No FT) and after LoRA fine-tuning on DreamBench++ and DiffusionDB. As shown in Tab. 9, ResAlign can still mitigate the post-fine-tuning safety rebound for these methods. Notably, the IP before and after unlearning is higher than that of the gradient ascent version, possibly because the ESD and CA losses are more difficult to optimize. We provide further discussions in Appendix E.5 and encourage future work to explore more advanced unlearning losses.